OpenStudy (anonymous):

What are the solutions of the quadratic equation? 2x^2-16x+32=0

OpenStudy (anonymous):

Also, I have to show my work

OpenStudy (anonymous):

Factor, you get 2(x^2 - 8x + 16) = 0 2(x-4)(x-4) = 0 set each factor equal to 0, you get x = 4.

OpenStudy (gorv):

2x^2-8x-8x+32=2x(x-4)-8(x-4)=(2x-4)(x-4) x=4 and 2

OpenStudy (anonymous):

 Okay, so fit those three values into the quadratic formula, which is;